OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4). y = -2x + 5 y = 2x - 3 y = 2x + 6 y = -2x + 2

OpenStudy (zale101):

use the equation y=mx+b where m=slope and b=y-intercept

OpenStudy (whpalmer4):

use the slope equation: \[m = \frac{y_2-y_1}{x_2-x_1}\] to find the slope of the line passing through \((x_1,y_1)\) and \((x_2,y_2)\)Armed with the slope, use the slope-point form to write the equation of a line with slope \(m\) passing through point \((x_0, y_0)\): \[y-y_0 = m(x-x_0)\] Then rearrange into slope-intercept form: \[y=mx+b\]